### v4

# Formal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents

Abstract
Recent advancements on Large Language Models (LLMs) enable AI Agents to automatically generate and execute multi-step plans to solve complex tasks. However, since LLM’s content generation process is hardly controllable, current LLM-based agents frequently generate invalid or non-executable plans, which jeopardizes the performance of the generated plans and corrupts users’ trust in LLM-based agents. In response, this paper proposes a novel “Formal-LLM” framework for LLM-based agents by integrating the expressiveness of natural language and the precision of formal language. Specifically, the framework allows agent developers to express their requirements or constraints for the planning process as an automaton. A stack-based LLM plan generation process is then conducted under the supervision of the automaton to ensure that the generated plan satisfies the constraints, making the planning process controllable. We conduct experiments on both benchmark tasks and practical real-life tasks, and our framework achieves over 50% overall performance increase, which validates the feasibility and effectiveness of employing Formal-LLM to guide the plan generation of agents, preventing the agents from generating invalid and unsuccessful plans. Further, more controllable LLM-based agents can facilitate the broader utilization of LLM in application scenarios where high validity of planning is essential. The source code of this work is available at https://github.com/agiresearch/Formal-LLM.
Formal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents
Zelong Li, Wenyue Hua, Hao Wang, He Zhu, Yongfeng Zhang Department of Computer Science, Rutgers University, New Brunswick {zelong.li, wenyue.hua, hw488, hz375, yongfeng.zhang}@rutgers.edu
1 Introduction
Numerous applications have emerged with the rapid development of Large Language Models (LLM). One notable application is the LLM-based agent, which is capable of automatically generating and executing multi-step plans to solve complex tasks. While the LLM-based agent exhibits creativity, there is a concern about the potential generation of unreasonable and invalid plans, undermining the effectiveness of agents. For example, generating a plan that attempts to process image data using a tool designed for text can lead to errors. Recent studies have pointed out the challenges of LLM-based agents in developing non-executable plans without sufficient human oversight (Ge et al., 2023a; Yuan et al., 2024). Addressing these challenges is crucial for improving agent performance, increasing the generation of valid plans, and maintaining user trust. Various attempts have been made to control LLM text generation, such as incorporating hard constraints (Takase and Okazaki, 2019; Carlsson et al., 2022), soft constraints (Gu et al., 2022; Lu et al., 2022), or a combination of both (Chen et al., 2024). However, the focus of controlling LLM-based agents emphasizes the validity of plans and tool use over purely text generation. Some studies (Ge et al., 2023a; Yuan et al., 2024) use LLM as a parser to extract a chain of tools from the generated texts based on prompts, yet satisfactory ratios for valid plans remain elusive.
To tackle the problems of invalid plan generation, we propose a framework named “Formal-LLM”, which integrates the expressiveness of natural language and the precision of formal language, as shown by the toy example in Figure 1. Specifically, to control the LLM-based agent’s plan generation, agent developers construct a context-free grammar (CFG) as the formal language to represent the constraints for the agent. The CFG is then automatically translated into a pushdown automaton (PDA). When LLM conducts planning, it is prompted to follow the state transition defined by the automaton. This is realized by limiting the LLM-based agent’s choices at each step to the valid actions defined by the PDA at its current state, which helps to guarantee that the constraint is satisfied in the final generated plan. We choose PDA in this work because some tasks need to be solved by tree-structured plans rather than chain-structured plans, and generating tree-structured plans requires a PDA.
Figure 1: The Formal-LLM workflow with a toy example. To control the LLM-based agent’s plan generation, agent developers construct a formal language (context-free grammar, CFG) to represent the natural language constraints. The formal language is then translated into a pushdown automaton (PDA). When LLM conducts planning, it needs to follow the state transition defined by the automaton, which helps to guarantee that the constraint is satisfied in the final generated plan. We choose PDA because some plans exhibit tree-structure (as shown by other examples in the paper), which requires a PDA to generate.
Furthermore, we introduce the backtracking mechanism in the Formal-LLM to increase the probability of finding valid plans for the agent, which enables the planning process to return to the previous step when a dead end is reached on the automaton. Besides, traditional LLM-based agent fine-tuning techniques such as Reinforcement Learning from Task Feedback (RLTF) (Ge et al., 2023a) rely on the rewards from the agent’s plan execution to fine-tune the LLM parameters. However, since many invalid plans may be generated, many rewards are actually not informative for LLM-based agent fine-tuning. Our Formal-LLM approach guarantees that invalid plans are excluded during the agent’s plan generation. As a result, our approach helps to increase the amount of valid rewards for LLM fine-tuning, which improves the performance of fine-tuned LLM-based agents.
In our experiments, we implement the Formal-LLM framework across various LLMs, including both open-source and closed-source LLMs. We also test our framework on both benchmark tasks and real-life practical tasks. Specifically, the benchmark tasks involve the LLM-based agent utilizing different tools to solve complex problems through multiple steps. The real-life scenarios encompass daily routines, cooking instructions, and commercial risk management, each imposing specific domain-knowledge or common-sense constraints. Our findings demonstrate the Formal-LLM’s capability to generate reasonable plans. The framework substantially enhances the overall performance on the benchmark tasks by over 50% and can always generate executable plans. In real-life scenarios, we provide qualitative analyses into the improvements brought by the Formal-LLM framework, affirming its feasibility and effectiveness in rendering the agents more controllable.
2 Related Work
2.1 LLM-based AI Agent
An AI agent is an autonomous entity capable of making decisions and executing actions within a specific environment to effectively tackle diverse complex tasks using AI techniques (Ge et al., 2023b; Mei et al., 2024; Wang et al., 2023; Xi et al., 2023). The emergence of Large Language Models (LLMs), exemplified by the GPT series (Radford et al., 2019; Brown et al., 2020; OpenAI, 2023) and the LLaMA series (Touvron et al., 2023a, b), has spurred the exploration of LLM-based agents (Ge et al., 2023a; Huang et al., 2022). These agents utilize LLMs as their central cognitive component or controller, broadening their perceptual and action capabilities through approaches like multimodal perception and tool usage (Schick et al., 2023; Ge et al., 2023a; Qin et al., 2023). In contrast to pre-LLM era AI agents, LLM-based agents showcase creativity, manifesting in the ability to generate innovative ideas without additional learning (Franceschelli and Musolesi, 2023), indicating a degree of self-directed exploration and decision-making (Xi et al., 2023). Despite LLM-based agent’s application in diverse real-world scenarios such as software development (Li et al., 2023a; Qian et al., 2023), scientific research (Boiko et al., 2023), and system management (Liu et al., 2023), recent studies have highlighted issues of non-executable plans being generated without adequate human oversight (Ge et al., 2023a; Yuan et al., 2024). If these plans lack executability, the utility of LLM-based agents in fields that require high validity is compromised, and their unreliability erodes user trust. To address this challenge, we propose integrating natural language and precise automaton during the LLM-based agent’s planning process.
2.2 Controllable LLM Generation
To the best of our knowledge, there are few works on controllable LLM-based agents, and discussions primarily revolve around controllable LLM text generation. Controlling LLM text generation typically falls into two categories: hard constraints and soft constraints (Qin et al., 2022). Hard constraints involve constraints on the desired text length, designated keywords for the generated text (Takase and Okazaki, 2019; Carlsson et al., 2022), and constraint vocabulary space in decoding time (Hemmer et al., 2023; Geng et al., 2024, 2023), while soft constraints restrict the output based on specific semantics, such as sentiments or topics (Gu et al., 2022; Lu et al., 2022; Li et al., 2020). A recent work (Chen et al., 2024) attempts to integrate both constraints into a unified approach. However, text generation control has disadvantages and problems. First, text generation control could hardly be directly applied to closed-source LLMs due to their need to access the decoding probability (Geng et al., 2024). Second, controllable LLM-based agents prioritize the validity of plan and tool use over purely text generation. Recent studies on LLM-based agents have noted limited effectiveness when employing text constraint generation on LLM-based agents and have explored using LLMs as a parser to extract planning information from generated texts (Ge et al., 2023a; Yuan et al., 2024); but relying heavily on the parser LLM’s effectiveness may not yield a satisfactory plan valid rate. We propose an automaton-guided agent planning to ensure that the generated plans adhere 100% to the constraints.
3 Preliminary Knowledge
3.1 Context-free Grammar
Many LLM agents operate tools that take multiple inputs and produce one output. For example, the tool “Visual Question Answering” receives an image and a text as the input and outputs a text. The plan involving such tools forms a tree structure, and such tree-structured plans require Context-Free Grammar (CFG) to express the rules.
The CFG consists of four components: terminals (symbols unable to be replaced), nonterminals (symbols subject to replacement), start symbol (a unique nonterminal, usually denoted as $S$), and productions (rules governing symbol substitution). The CFG production format follows:
$A\rightarrow\alpha$ (1)
where $A$ is a single nonterminal, and $\alpha$ is a string comprising any combination of terminals and nonterminals. $A$ can be replaced by $\alpha$ in any situation.
$\begin{split}&S\rightarrow\varepsilon|aSb\end{split}$ (2)
Eq.(2) is a CFG example with two productions separated by “||||”, where $a,b$ are terminals, $S$ is the start symbol and the only nonterminal, and $\varepsilon$ stands for the empty string. From the start symbol, we can use these two productions to generate infinite strings, such as the empty string, $ab$, and $aabb$. Take the derivation of $aabb$ as an example. We first apply the second production twice ($S\rightarrow aSb\rightarrow aaSbb$) and then the first production once ($aaSbb\rightarrow aabb)$. The formal language, Context-Free Language (CFL), contains the set of strings of terminals, which can be derived from $S$ in a certain amount of CFG production steps.
Figure 2: A Pushdown Automaton (PDA) example.
Figure 3: An equivalent PDA to the CFG of the combination of Eq.(3) and Eq.(4). The lowercase letters $a$ to $f$ are $a_{i}$ to $f_{j}$ in Eq.(3). For example, the transit $(b,B;\varepsilon)$ implies the PDA removes the stack top $B$ and remains in state $q_{1}$ if the PDA is at state $q_{1}$, the stack top is $B$, and the next string letter is $b_{1}/b_{2}/b_{3}$.
3.2 Pushdown Automaton
Every CFG can be transformed into an equivalent nondeterministic Pushdown Automaton (PDA) with an algorithm (Hopcroft et al., 2001). A PDA is a machine that moves through a series of states based on a given string input, a transition function, and the top element of a stack. The automaton reads the letters in the input sequentially, and if there exists a path leading the machine to an accept state right after consuming the entire string, the input string is considered an accepting word. The set of all accepting words is referred to the CFL accepted by the automaton. Since the PDA process is more intuitive and automated than CFG for guiding the prompting of LLMs, our framework is designed on top of PDA.
Figure 2 shows how PDA works as an example. This automaton includes a stack alphabet ($\{S,Z\}$) with an initial stack symbol $Z$, a start state $q_{0}$, an accept state set $\{q_{2}\}$, and a transit function over states. The transit functions are in the form of triplets. For example, the edge $(a,Z;SZ)$ from $q_{0}$ to $q_{0}$ means that if the stack top is $Z$ and the next letter of the string is $a$, the PDA remains in $q_{0}$ state and the stack top changes from $Z$ to $SZ$ ($S$ is the stack top now).
This PDA is equivalent to the CFG in Eq.(2), i.e., every accepting word of this PDA belongs to the CFL defined in Eq.(2). We still use $aabb$ as an example to show the PDA mechanism. Initially, the PDA is at $q_{0}$ and the stack only contains $Z$. We first apply $(a,Z;SZ)$ when receiving the first letter $a$, and the PDA remains in $q_{0}$ with the stack as $SZ$. Then, when consuming the next letter $a$, we use $(a,S;SS)$ and $(\varepsilon,S;S)$, move to $q_{1}$ state, and change the stack to $SSZ$. Next, the PDA remains in $q_{1}$ and applies $(b,S;\varepsilon)$ function twice to consume the two $b$’s in the input string, and at the same time, the stack transforms from $SSZ$ to $Z$. Finally, the automaton use $(\varepsilon,Z;Z)$ to move to the accept state $q_{2}$. Since the machine finishes reading the input string, $aabb$ is an accepting word of this CFL. $aabbb$ is not an accepting word because, after reading $aabb$, the automaton is at state $q_{1}$ with the stack as $Z$, and there is no valid function to consume the final $b$ with $Z$ as the stack top.
4 The Formal-LLM Framework
4.1 Motivation and Challenge
Natural language is easily understandable by humans, but it may also lack precision in certain application scenarios. In contrast, formal language is defined in clear mathematical and machine readable forms, which could be hard to comprehend for humans, but it also possesses great precision. Thus, we aim to effectively integrate the benefits of both natural and formal languages through the pushdown automaton (PDA) for more controllable planning of LLM-based agents in scenarios requiring precise and valid plans.
Constructing our Formal-LLM framework presents non-trivial challenges. First, LLMs may struggle to directly comprehend and process formal language due to the limited CFG- or PDA-related corpus in pre-training data. Hence, natural language prompts are necessary to effectively describe the status of the automaton, bridging the gap between formal and natural language for LLM’s planning. Given that the automaton has the potential to generate an infinite number of plans, the second challenge arises because the automaton only ensures the validity and executability of the plans by adhering to defined constraints, but cannot guarantee the performance and optimality of the generated plan. Therefore, leveraging the solid natural language understanding ability of LLM and other methods, such as fine-tuning and backtracking mechanisms, becomes crucial for enhancing planning performance.
4.2 Formulating Constraint to Automaton
The overall flow of our framework is presented in Figure 1. We begin by illustrating the process of converting constraints into formal language and automata for benchmark tasks, using the OpenAGI benchmark (Ge et al., 2023a) as an example, which incorporates a set of domain expert models as tools and lists a series of intricate problems that cannot be addressed with a single tool. These tools are categorized into six primary groups according to the input and output modalities, as indicated in Table 3 in the Appendix.
If the plans for the benchmark tasks generated by the LLM-based agent deviate from the expected data format, the execution performance degrades significantly or even results in errors. For instance, the Colorization tool operates on image data as input and output. Since image data is represented as a 3-D tensor, providing a 1-D tensor text string as input could lead to errors, due to incorrect data dimensions. Thus, the entire plan becomes invalid and cannot be executed. In such cases, natural language may struggle to articulate these constraints precisely and concisely, but formal language can express them straightforwardly.
The CFG in Eq.(3) outlines the constraints on data formats with Polish prefix notation: $T$ for text and $I$ for image, capital letters $A$ to $F$ for the six types of tools, and lowercase letters such as $a_{i}\in A$ or $f_{j}\in F$ for specific tools in the order of Table 3.
$I\rightarrow AI|CT\qquad$ $T\rightarrow BI|DT|EIT|FTT$ (3)
$A\rightarrow a_{1}|a_{2}|a_{3}|a_{4}\qquad$ $B\rightarrow b_{1}|b_{2}|b_{3}$
$C\rightarrow c_{1}\qquad$ $D\rightarrow d_{1}|d_{2}|d_{3}|d_{4}|d_{5}$
$E\rightarrow e_{1}\qquad$ $F\rightarrow f_{1}$
where $I\rightarrow AI|CT$ means $I$ can be replaced by $AI$ or $CT$, because $A$ is an image-to-image tool and $C$ is a text-to-image tool. As a result, when $A$ is applied on an image ($AI$) or $C$ is applied on a text ($CT$), the result is still an image. Considering a specific task, such as “Given blurry gray-scale images, how to return the object names in English step by step?”, we know that the task’s inputs are images and the final outputs should be texts. To formalize this, we add two constraints to Eq.(3):
$S\rightarrow T\qquad$ $I\rightarrow i$ (4)
i.e., the CFG starts with the symbol $T$, implying the final text format, and uses the lowercase letter $i$ to represent the input image. The combination of Eq.(3) and Eq.(4) renders the data format constraints of the specific task. For instance, $e_{1}a_{1}ib_{1}i$ is an accepted word, representing a valid plan: “Utilize Image Classification ($b_{1}$) on the input image to obtain text and apply Colorization ($a_{1}$) on the input image to generate the output image. Then, employ Visual Question Answering ($e_{1}$) with the text and output image to derive the final text.” As mentioned in Section 3.2, the CFG can be equivalently converted into a PDA in Figure 3. These steps allow us to transform the constraints into automata for subsequent utilization.
In certain scenarios, creating an automaton can be more straightforward than formulating a formal language. Section A.5.1 describes a real-life task for daily work planning. The constraints in this case revolve around time, allowing us to represent time as distinct states, as depicted in Figure 4. The automaton initiates the design of an activity that concludes at 20:00, and a valid plan is generated when the automaton reaches 10:00 while being in a state where three meals have been consumed. In this case, we design a PDA without a CFG.
4.3 Formal-LLM Prompts and Planning from Automaton
In Section 4.2, we convert the natural language planning constraints into a PDA, where any accepted word represents a valid and executable plan. The reason is that LLM could struggle to directly comprehend or process the CFG due to limited exposure to formal language during pre-training. Hence, we employ natural language prompts for LLM to comprehend the task and generate plans that can be easily read by humans; simultaneously, we use the PDA to guide the process of generating natural language plans.
The automaton initiates the plan generation process from the start state. In situations where multiple viable options are available for the automaton to proceed, a prompt is created to inquire the LLM, as shown in Figure 1. This prompt encompasses a comprehensive {task description} to mitigate potential forgetting issues of the LLM (Hua et al., 2023), details about the {current progress}, and {feasible choices} determined by the PDA. For example, when the PDA in Figure 3 is in state $q_{1}$ with the symbol $I$ at the top of the stack, three viable transits: $(\varepsilon,I;AI)$, $(\varepsilon,I;CT)$, and $(i,I;\varepsilon)$ are included in the {feasible choices} of the prompt. A prompt example is provided in Appendix A.3.
Another advantage of employing PDA for LLM-based agent planning is its ability to articulate plans that involve tools with many-input-single-output (e.g., Image-Text pair as input, Text as output), assuming the number of inputs of each tool is deterministic. To illustrate, we depict the derivation tree of the string $e_{1}a_{1}ib_{1}i$ in Figure 5. Treating the tools $a_{1}$ and $b_{1}$ as functions, the string is functionally equivalent to $t=e_{1}\big{(}a_{1}(i),b_{1}(i)\big{)}$. When the PDA finishes the substitution of symbols $E$ and $I$, it can well know that the stack symbol $T$ is derived from $(\varepsilon,T;EIT)$ and matches the input of tool $e_{1}$, because they are pushed into the stack at the same time. Finally, we can identify the tools that will operate on this image/text data in the subsequent step and incorporate this information in the prompt, enabling the LLM to make informed decisions.
Figure 4: The PDA includes 13 states, $\{10,11,...,19,20,Start,Fin\}$, with $Start$ as the start state and $\{Fin\}$ as the accept state set. The lowercase letter $a$ is for breakfast, $b$ is for lunch, $c$ is for supper, $d$ is for playing basketball, $e$ is for grocery shopping, $f$ is for house cleaning, $g$ is for homework, and $h$ is for laundry. The stack symbols $D,Z,E,N$ represent 3, 2, 1, and 0 meals remaining to plan, respectively. $\forall$ represents any stack symbol. The variable $i$ is enumerated from 20 to 11, $j$ is enumerated from 20 to 11 except 13, and $k$ is enumerated from 20 to 13.
Figure 5: The derivation tree for $e_{1}a_{1}ib_{1}i$ for the example benchmark task in Section 4.2.
4.4 Reinforcement Learning from Task Feedback and Backtracking Mechanism
An accepting word of an automaton can be infinitely long. However, we anticipate that the plan proposed by the LLM-based agent can be executed within a finite number of steps. Consider the PDA in Figure 2, any length of the string in the form of $a^{n}b^{n}$ where $n\geq 0$ is an accepting word. Similarly, in planning tasks, tools can be employed multiple times, such as applying the Text Summarization tool to the original text and infinitely repeating the same tool on the output. However, such planning behavior may be nonsensical and inefficient. Hence, for the tasks outlined in this paper, each tool is restricted to a single use for a given task, preventing infinite planning for both benchmark and real-life scenarios.
Upon imposing constraints on the tool usage, a new challenge emerges: the automaton may encounter a dead-end during plan generation. Consider the formal language in Eq.(3) and Eq.(4). Upon analyzing the language, we observe that the only means of transitioning from the symbol $T$ to $I$ (and subsequently to a terminal $i$) involve $B$-type (Image-in, Text-out) and $E$-type (Image-Text-pair-in, Text-out) tools. If these two tool types are exhausted, yet there are still $T$ symbols in the stack, the automaton will inevitably fail to reach the accept state $q_{2}$, as no methods remain to clear the symbol $T$. Finally, a dead-end situation may arise following the imposition of tool usage limits.
To address this challenge, we propose to document the automaton’s details at each step, encompassing the current state and stack, and the part of the word that has been generated. When the automaton confronts a dead-end, we initiate backtracking to the preceding steps until there is an unexplored branch. The backtracking involves retracing the automaton to the prior step, reinstating the automaton details based on the recorded stack, and eliminating the dead-end branch from the prior choice list. Thus, the automaton is capable and guaranteed to generate a valid plan if it exists.
The design above only guarantees the validity of the generated plans, but not the optimal plan. Thus, as an additional enhancement in plan quality, particularly when employing open-source LLMs, we integrate Reinforcement Learning from Task Feedback (RLTF) (Ge et al., 2023a) following the application of our Formal-LLM framework. After the LLM-based agent generates a plan for a training task, the plan is executed on benchmark data to assess its performance. Then, the performance is used as a reward for RL to update the parameters of the LLM. Our framework ensures that invalid plans are always excluded during the LLM-based agent’s plan generation. Therefore, the number of valid rewards for LLM fine-tuning increases, and our approach can improve the performance of fine-tuned LLM-based agents.
5 Experiments
5.1 Backbone Large Language Model (LLM)
We test our Formal-LLM framework on these closed-source LLMs:
  • GPT-3.5-turbo (Brown et al., 2020) is a generative pre-trained transformer of OpenAI.
  • Claude-2 (Claude-2, 2023) is a transformer LLM of Anthropic.
  • GPT-4 (OpenAI, 2023) is a follow-up version of GPT-3.5.
and these open-source LLMs:
  • Flan-T5-Large (Chung et al., 2022) is a language model with 770 million parameters.
  • Vicuna-7B (Chiang et al., 2023) is a 7-billion size chatbot trained by fine-tuning the LLaMA model (Touvron et al., 2023a).
  • LLaMA-2-13B (Touvron et al., 2023b) is a successor to the 13-billion size LLaMA model.
5.2 Learning Schema of LLMs
We adopt the following LLM learning schema:
  • Zero-shot Learning (Zero) directly inputs the prompt to the LLM.
  • Few-shot Learning (Few) presents a set of high-quality examples in the prompt, each consisting of input and desired output on the target task.
  • Reinforcement Learning from Task Feedback (RLTF) applies text constraint generation, executes the plan, and takes its performance as the reward to optimize the LLM by RL.
  • Formal-LLM (F-LLM) is our proposed framework, using automaton to control planning.
  • Formal-LLM plus RLTF (F-LLM+RLTF) applies our Formal-LLM framework on top of RLTF without text constraint generation. We use automaton to exclude invalid plans.
Specifically, we utilize Zero, Few, and F-LLM frameworks for closed-source LLMs, as these frameworks do not need to modify the LLM’s parameters. Regarding open-source LLMs, we compare RLTF, F-LLM and F-LLM+RLTF, given that RLTF surpasses Zero and Few, as exhaustively discussed in recent work (Ge et al., 2023a).
Metrics / Task GPT-3.5-turbo Claude-2 GPT-4
Zero Few F-LLM (Ours) Zero Few F-LLM (Ours) Zero Few F-LLM (Ours)
% of Valid Plans 29% 71% 100% 29% 47% 100% 53% 76% 100%
Task 1 (CLIP Score) 0.0 0.0 0.3056 0.0 0.2543 0.3056 0.0 0.3055 0.3056
Task 2 (BERT Score) 0.1914 0.3820 0.6364 0.2111 0.5038 0.6275 0.2076 0.6307 0.5102
Task 3 (ViT Score) 0.2437 0.7497 0.6470 0.4082 0.5416 0.7137 0.5058 0.6480 0.7689
Task X 0.0 0.0 0.0658 0.0 0.0 0.2799 0.0 0.0 0.2876
Average over tasks 0.1443 0.3345 0.4846 0.1838 0.3773 0.5420 0.1992 0.4662 0.4914
Table 1: Benchmark task performances under different settings for three closed-source LLMs. Zero is for Zero-shot Learning, Few is for Few-shot Learning, and F-LLM is for Formal-LLM. The boldface numbers denote the highest score under each task type using the same LLM.
Metrics / Tasks Flan-T5-Large Vicuna-7B LLaMA-2-13B
RLTF F-LLM F-LLM+RLTF RLTF F-LLM F-LLM+RLTF RLTF F-LLM F-LLM+RLTF
% of Valid Plans 24% 100% 100% 29% 100% 100% 47% 100% 100%
Task 1 (CLIP Score) 0.0 0.3049 0.3049 0.0 0.3122 0.3139 0.0610 0.1601 0.3060
Task 2 (BERT Score) 0.3327 0.5164 0.5287 0.1475 0.4948 0.4673 0.1611 0.4220 0.5565
Task 3 (ViT Score) 0.6632 0.6264 0.7469 0.6958 0.5948 0.8618 0.7106 0.7043 0.6808
Task X 0.0 0.0728 0.4046 0.0 0.4127 0.4029 0.0 0.3846 0.4163
Average over tasks 0.3111 0.4451 0.5321 0.2009 0.4824 0.5162 0.3101 0.4498 0.5390
Table 2: Benchmark task performances under different settings for three open-source LLMs. RLTF is for Reinforcement Learning from Task Feedback, F-LLM is for Formal-LLM, and F-LLM+RLTF is for using the generated plans by F-LLM to calculate the reward for RLTF. The boldface numbers denote the highest score under each task type using the same LLM.
5.3 Experimentation Datasets
5.3.1 Benchmark Datasets
We conduct experiments on two benchmark datasets, i.e., OpenAGI (Ge et al., 2023a) and TravelPlanner (Xie et al., 2024). Due to space limit, we provide the results on OpenAGI in the following and provide the results on TravelPlanner in Appendix Section A.4. The OpenAGI benchmark tasks are categorized based on their output type and ground-truth label type (Task 1, 2, and 3). Then, based on different task types, different metrics are employed to gauge the performance: CLIP Score (Hessel et al., 2021), assessing the similarity between text and image, is utilized for Text-to-Image tasks; BERT Score (Zhang et al., 2020), evaluating text generation with BERT, is applied when both data labels and the expected outputs are texts; and ViT Score (Wu et al., 2020) gauges the similarity between the image label and image output. Additionally, we construct Task X, which is a subset of “Task 1 $\cup$ Task 2 $\cup$ Task 3” that requires a tree-structured plan due to the use of many-input-single-ouput tools, such as Question Answering. Task X is used to test the complex planning ability of our Formal-LLM framework, and is evaluated using the corresponding metric. The tools are detailed in Table 3 and example tasks of each category are in Table 4 in the Appendix.
5.3.2 Real-life Practical Tasks
We also experiment with real-life planning scenarios, encompassing daily plans, cooking recipes, and risk management, where validity and rationality are crucial. In these scenarios, the concept of tool is generalized to include various types of steps in a plan, such as events, actions, or activities, as they can take diverse forms supporting the execution of the plan. For example, a tool in the daily planning task may be an activity such as taking breakfast. We provide qualitative analyses between the Zero and F-LLM learning schema using GPT-4 backbone for practical tasks. We use GPT-4 for the experiment of practical tasks because other LLMs can hardly generate readable plans for these tasks, and we test under Zero and F-LLM due to the limited sample size and that these two learning frameworks do not require accessing to the LLM parameters.
5.4 Experimental Analysis
The experiment results on the benchmark tasks are shown in Table 1 and Table 2, referring to the closed-source and open-source LLMs, respectively. Each row stands for a type of task, each column represents the learning schema of an LLM-based agent, and every three columns are the results of the same LLM. For closed-source LLMs, among the three learning schemata without revising the LLM parameter, almost all the best scores under each type of task belong to our F-LLM framework. For open-source LLMs, our F-LLM without fine-tuning is already better than RLTF (the best schema in the OpenAGI platform (Ge et al., 2023a)) in most cases, except for the ViT Score (Task 3), because the Type-3 tasks take the most significant portion of the RLTF’s fine-tuning data, and thus RLTF is adequately optimized on these tasks. The performance gain of our F-LLM framework comes from the 100% executable plans and supporting tree-structured planning for the difficult tasks in Task X. As a comparison, the best open-source LLM (LLaMA-2-13B with RLTF) can only generate 47% executable plans, and the best closed-source LLM (GPT-4 with few shot) can generate 76% executable plans. Additionally, none of the baselines can handle the Task X (Score = 0.0). Due to the increased amounts of valid rewards from the 100% executable plans with our framework, the F-LLM+RLTF approach enables boosted performance, as demonstrated by its better scores in Table 2. To summarize, the benchmark experiment results show that our Formal-LLM framework is an effective method to integrate the benefits of both natural and formal languages for more controllable and valid planning of LLM-based agents.
5.5 Case Study
We show the complete results of real-life examples by applying our Formal-LLM to the GPT models in the Appendix. From the results, after applying our Formal-LLM, the generated plan from the GPT-based agent is more complete, reasonable and specific to the case. In the daily plan example, the agent fails to fit all the activities into the plan without the Formal-LLM, while it can achieve this goal by applying the Formal-LLM. Still take Figure 4 as an example, our Formal-LLM can limit the planning to 10:00$\sim$20:00, while without Formal-LLM, the agent schedules activities after 20:00, even if we mention the constraint “Generate a plan for all the activities between 10:00 and 20:00” in natural language, which shows the advantages of formal language guided planning. For the cooking recipe example, without strict automaton constraint, GPT may generate steps such as “stir-fry the Chinese broccoli until it changes color” after “once the water boils, add the Chinese broccoli”. This is unreasonable because broccoli cannot be stir-fried in boiling water. For the risk management example, the generated plan without our framework is too general that it could be used for any other two companies. However, the plan after applying Formal-LLM is more specific and focuses on the potential antitrust risk between Microsoft and Blizzard. Thus, the Formal-LLM can use the automaton describing the constraints to generate higher-quality plans.
6 Conclusions and Future Work
In this study, we introduce the innovative Formal-LLM framework for LLM-based agents, combining LLM’s power of natural language comprehension with the precision of formal language. Our experiments, encompassing benchmark tasks and real-life practical scenarios, affirm the feasibility and effectiveness of employing automaton to control the agent’s generation of valid plans. More controllable LLM-based agents can augment the potential utilization of LLM in applications where high validity of planning is important.
Our work has several potential extensions. First, automating the translation of natural language into formal language for agents could further improve the framework.Additionally, this work focuses on LLM plan generation based on formal language. Another important problem to explore is LLM plan verification based on formal language.
Acknowledgment
We thank Harun Taha Kepenek for providing helpful suggestions to improve this research paper.
Appendix A Appendix
A.1 OpenAGI Benchmark Tasks and Tools
The tool list integrated in the OpenAGI platform (Ge et al., 2023a) is shown in Table 3 and the different types of tasks as well as example tasks are shown in Table 4.
Input Modality Output Modality Tool Description Expert Model
Image Image ($A$) Colorization ($a_{1}$) ColorizerFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Zhang et al., 2017)
Super-Resolution ($a_{2}$) Swin2SRFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Conde et al., 2022)
Image Denoising ($a_{3}$) RestormerFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Zamir et al., 2022)
Image Deblurring ($a_{4}$) RestormerFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Zamir et al., 2022)
Text ($B$) Image Classification ($b_{1}$) ViTFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Dosovitskiy et al., 2021)
Object Detection ($b_{2}$) DETRFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Carion et al., 2020)
Image Captioning ($b_{3}$) Vision Encoder DecoderFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Li et al., 2023b)
Text Image ($C$) Text-to-Image Generation ($c_{1}$) Stable DiffusionFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Rombach et al., 2022)
Text ($D$) Sentiment Analysis ($d_{1}$) DistilBERTFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Sanh et al., 2020)
Text Summarization ($d_{2}$) BARTFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Lewis et al., 2019)
Machine Translation ($d_{3}$) T5Formal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Raffel et al., 2020)
Fill Mask ($d_{4}$) DistilBERTFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Sanh et al., 2020)
Text Generation ($d_{5}$) GPT-2Formal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Radford et al., 2019)
Image-Text Pair Text ($E$) Visual Question Answering ($e_{1}$) GITFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Wang et al., 2022)
Text-Text Pair Text ($F$) Question Answering ($f_{1}$) DistilBERTFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Sanh et al., 2020)
Table 3: Tool list integrated in OpenAGI platform (Ge et al., 2023a) for benchmark tasks. The tools are categorized into six primary groups according to the input and output modalities.
A.2 Implementation Details
Our framework and all baselines are implemented by PyTorch, an open-source library. We follow the implementation setting of the OpenAGI platform (Ge et al., 2023a) for baselines. For the F-LLM+RLTF framework, we use the REINFORCE (Williams, 1992) as the core reinforcement learning algorithm of RLTF. We use the original checkpoint of each backbone LLM without supervised fine-tuning. We set the maximum number of updating epochs at 30 and use Adam as the optimizer with the learning rate at 0.001 for RLTF. Also, we apply Low-Rank Adaptation (LoRA) (Hu et al., 2021) to the RLTF for efficient fine-tuning with the rank as 8 and the placement as $v_{proj}$ ($W_{v}$) and $q_{proj}$ ($W_{q}$). Our experimental results are the average of five runs.
A.3 Prompt for OpenAGI Benchmark Tasks
An example of the CFG and the corresponding PDA are shown in Eq.(3) and Eq.(4) as well as Figure 3. Different benchmark tasks share a common subset of constraints shown in Eq.(3) corresponding to the input-output constraints of the tools. Besides, each task has its own constraint on the input-output of the task, resulting in different constraints in Eq.(4), and thus the final PDA (Figure 3) for different tasks are different. Also, we directly use the tool name (e.g., Colorization) instead of the category name (e.g., “Image-in, Image-out tools”) in the prompt. For example, consider the PDA in Figure 3 when it is in state $q_{1}$ with the symbol $I$ at the top of the stack. There are three viable transits: $(\varepsilon,I;AI)$, $(\varepsilon,I;CT)$, and $(i,I;\varepsilon)$. In practice, as nonterminals $A$ and $C$ are promptly replaced by $terminals$ from $a_{1}\sim a_{4}$ and $c_{1}$. This is because the tool name provides more concrete information for the LLM to comprehend the tool’s functionality, enabling the LLM to make informed decisions without influencing the plan’s executability.
Zero-shot Prompt:
Problem: {task_description}.
What is its soltuion? Use Setp to mark.
Few-shot Prompt:
{few shot examples in the format of:
Problem: {task_description}.
Solution:
Step 1: ...
Step 2: ...
Step k: ...
Problem: {task_description}.
Solution:
RLTF Prompt (RLTF executes the solution and use the performance as reward to fine-tune the LLM):
Problem: {task_description}.
Solution:
Formal-LLM Prompt (zero-shot):
You will help me generate a plan for the problem: {task_description} by answering a series of my questions.
{current_progress}
To get the {data_modality}, we have the following choices:
{choice_list}
Your answer should be only an integer, referring to the desired choice.
Formal-LLM + RLTF (execute the solution of Formal-LLM and use the performance as reward to fine-tune the LLM):
You will help me generate a plan for the problem: {task_description} by answering a series of my questions.
{current_progress}
To get the {data_modality}, we have the following choices:
{choice_list}
Your answer should be only an integer, referring to the desired choice.
Formal-LLM Prompt Example:
You will help me generate a plan for the problem: "Given a grayscale image, how to return the regular image step by step?" by answering a series of my questions.
Current Progress:
Step n: Use Image Super Resolution;
Step (n-1): ?
To get the input image of "Image Super Resolution", we have the following choices:
1: the output of Colorization,
2: the output of Image Denoising,
3: the output of Image Deblurring,
4: the output of Text to Image Generation,
5: Input Image.
Your answer should be only an integer, referring to the desired choice.
Task Metrics Output Label Evaluation Task Example
Task 1 CLIP Score Image Text Text-to-Image similarity Given clozed English text,
how to generate an image
step by step?
Task 2 BERT Score Text Text Text-to-Text similarity Given noisy grayscale image,
how to return the caption in
German step by step?
Task 3 ViT Score Image Image Image-to-Image similarity Given blurry grayscale image,
how to return the regular
image step by step?
Task X Corresponding Score / / / Given low-resolution noisy
blurry grayscale image and
English query, how to
answer the question in
German step by step?
Table 4: Benchmark task examples under each category. Task X is a subset of “Task 1 $\cup$ Task 2 $\cup$ Task 3”, which is a subset of tasks that require a tree-structured plan rather than chain-structured plan due to the use of many input-single-output tools. Task X is used to test the complex planning ability of our Formal-LLM framework.
A.4 Formal-LLM on TravelPlanner Benchmark
We apply our Formal-LLM framework to the TravelPlanner (Xie et al., 2024) platform. TravelPlanner is a benchmark designed to assess language agents’ performance in tool use and complex planning under various constraints. We use GPT-4 as the backbone LLM because TravelPlanner’s experiments show that other LLMs cannot generate valid plans that meet all constraints. We use the workflow in Figure 6 as the automaton to control the agent planning in the Formal-LLM framework. The flowchart is essentially a Deterministic Finite Automaton (DFA), which is a special case of Pushdown Automaton (PDA). We provide an introduction to the experiment in the following.
A.4.1 Evaluation Metrics
  • Delivery Rate: This metric assesses whether agents can successfully deliver a final plan.
  • Commonsense Constraint Pass Rate: This metric evaluates whether a plan satisfies the pre-defined eight commonsense constraints.
  • Hard Constraint Pass Rate: This metric measures whether a plan meets all explicit constraints in the given query.
  • Final Pass Rate: This metric represents the proportion of the plans that meet all aforementioned constraints among all tested plans.
Besides, for the Commonsense Constraint Pass Rate and Hard Constraint Pass Rate, we report the Micro Pass Rate, calculating the ratio of satisfied constraints to the total number of constraints, and the Macro Pass Rate, calculating the ratio of plans that meet all commonsense and hard constraints among all tested plans.
A.4.2 Baseline
We utilize GPT-4 as the backbone LLM and adopt the ReAct (Yao et al., 2023) framework as our baseline setting, as it demonstrated the highest performance in previous experiments (Xie et al., 2024).
Figure 6: The flowchart for TravelPlanner benchmark.
Planning Strategy Delivery Rate Commonsense Pass Rate Hard Constraint Pass Rate Final Pass Rate
Micro Macro Micro Macro
ReAct (Yao et al., 2023) 93.1 63.3 2.0 10.5 5.5 0.6
Formal-LLM (Ours) $100$ $82.9$ $16.7$ $41.3$ $31.7$ $9.2$
Table 5: Experiment results on the testing dataset of TravelPlanner (Xie et al., 2024) benchmark when using GPT-4 as the backbone LLM. The best results are marked in bold.
A.4.3 Experiment Result
The experiment results on the TravelPlanner benchmark are shown in Table 5. Each row stands for a planning strategy and each column represents an evaluation metric. The results show that our Formal-LLM is better than the best baseline in every metric, validating that our Formal-LLM can generate more reasonable plans. Also, unlike the experiments on the OpenAGI platform, Formal-LLM does not generate 100% valid plans on the TravelPlanner benchmark, due to the task complexity and automaton quality. However, the generated plans meet the constraints described in the automaton in Figure 6. We can improve the automaton quality for TravelPlanner and report even better performances in the future.
A.5 Prompt for Real-life Tasks
A.5.1 Daily Plan
The automaton is displayed in Figure 4.
GPT-4 Prompt:
Generate a plan for activities between 10:00 and 20:00.
Breakfast, lunch, and supper need 1 hour each.
Outdoor activities: basketball playing 13:00 - 15:00; do grocery shopping needs 1 hour.
Indoor activities: house cleaning needs 1 hour; homework needs three hours; turning on the washer/laundry machine needs 0 minutes but needs to stay home for one hour.
Other constrain:
Cannot play basketball within an hour after a meal.
Formal-LLM Prompt:
Generate a plan for activities between 10:00 and 20:00.
Breakfast, lunch, and supper need 1 hour each.
Outdoor activities: basketball playing 13:00 - 15:00; do grocery shopping needs 1 hour.
Indoor activities: house cleaning needs 1 hour; homework needs three hours; turning on the washer/laundry machine needs 0 minutes but needs to stay home for one hour.
Other constrain:
Cannot play basketball within an hour after a meal.
Lets start planning from the end.
{current_progress}
Decide on the activity ending at {current_hour}:00.
Here are possible options:
{choice_list}
Your reply should be only one number, such as 1, referring to the option.
Formal-LLM Prompt Example:
Generate a plan for activities between 10:00 and 20:00.
Breakfast, lunch, and supper need 1 hour each.
Outdoor activities: basketball playing 13:00 - 15:00; do grocery shopping needs 1 hour.
Indoor activities: house cleaning needs 1 hour; homework needs three hours; turning on the washer/laundry machine needs 0 minutes but needs to stay home for one hour.
Other constrain:
Cannot play basketball within an hour after a meal.
Lets start planning from the end.
Current Progress:
17:00 - 20:00 Doing homework.
Decide on the activity ending at 17:00.
Here are possible options:
1. Eating supper.
2. Grocery shopping.
3. House cleaning.
4. Turning on the washer/laundry machine.
5. Doing nothing for one hour.
Your reply should be only one number, such as 1, referring to the option.
A.5.2 Cooking Recipe
The formal language for the cooking task is a CFG, shown as Eq.(5).
$\begin{split}&S\rightarrow aAB\\ &B\rightarrow aCD|aCE\\ &C\rightarrow aF\\ &D\rightarrow aAI\\ &E\rightarrow bH\\ &F\rightarrow bAG\\ &G\rightarrow cd\\ &H\rightarrow ef|eI\\ &I\rightarrow cf\end{split}$ (5)
[…]
GPT-4 Prompt:
Generate a broccoli beef cooking plan.
The ingredients include raw beef slices, carpaccio, broccoli, onions, ginger, garlic, and water.
The seasonings include cooking oil, salt, light soy sauce, cooking wine, white pepper, sugar, vinegar.
Cooking utensils including woks and cooking pots.
Tableware including chopsticks, spoons, wok spoons, and several bowls.
Formal-LLM Prompt:
Generate a broccoli beef cooking plan.
The ingredients include raw beef slices, carpaccio, broccoli, onions, ginger, garlic, and water.
The seasonings include cooking oil, salt, light soy sauce, cooking wine, white pepper, sugar, vinegar.
Cooking utensils including woks and cooking pots.
Tableware including chopsticks, spoons, wok spoons, and several bowls.
{current_progress}
Decide on the previous step before current progress.
Here are possible options to get {target_item} for the step: {parent_step}.
{choice_list}
Your reply should be only one number, such as 1, referring to the option.
Formal-LLM Prompt Example:
Generate a broccoli beef cooking plan.
The ingredients include raw beef slices, carpaccio, broccoli, onions, ginger, garlic, and water.
The seasonings include cooking oil, salt, light soy sauce, cooking wine, white pepper, sugar, vinegar.
Cooking utensils including woks and cooking pots.
Tableware including chopsticks, spoons, wok spoons, and several bowls.
Current Progress:
Step n: Then, we get the cooked broccoli beef.
Step n-1: Stir-fry the beef and broccoli mixture with the seasoning in a wok.
Step n-2: Prepare the seasoning: ginger, garlic, cooking oil, salt, light soy sauce, cooking wine, white pepper for the step: "Stir-fry the beef and broccoli mixture with the seasoning in a wok."
Step n-3: ?
Decide on the previous step before current progress.
Here are possible options to get the mixture of beef and broccoli for the step: "Stir-fry the beef and broccoli mixture with the seasoning in a wok."
1: Combine lightly cooked beef and lightly cooked broccoli in a wok.
2: Combine lightly cooked beef and blanched and drained broccoli in a wok.
Your reply should be only one number, such as 1, referring to the option.
A.5.3 Risk Management
Figure 7: The flowchart for risk management task.
Domain experts draw a flowchart of risk management as Figure 7. The flowchart is essentially a Deterministic Finite Automaton (DFA), which is a special case of Pushdown Automaton (PDA). A question is generated to inquiry LLM when the automaton reaches a new state. The LLM can select a reasonable branch according to its knowledge of the specific seller, buyer, and company name. In the experiment, the seller is Shareholders of Blizzard Entertainment, the company is Blizzard Entertainment, and the buyer is Microsoft.
GPT-4 Prompt:
Task: You are a plan maker to design a risk management plan for deals related to companies.
Scenario: {Seller A} wants to sell the {Company B} which {Buyer C} wants to buy. Design a risk management plan for {Buyer C} before negotiating a detailed contract.
Formal-LLM Prompt:
Task: You are a plan maker to design a risk management plan for deals related to companies.
Scenario: {Seller A} wants to sell the {Company B} which {Buyer C} wants to buy. Design a risk management plan for {Buyer C} before negotiating a detailed contract.
{question in the flowchart}
{choice_list}
Your answer should be only one number, such as 1, referring to the option.
Formal-LLM Prompt Example:
Task: You are a plan maker to design a risk management plan for deals related to companies.
Scenario: Shareholders of Blizzard Entertainment wants to sell the Blizzard Entertainment which Microsoft wants to buy. Design a risk management plan for Microsoft before negotiating a detailed contract.
Question: Does Microsoft buying Blizzard Entertainment trigger mandatory antitrust filing?
1: Yes.
2: No.
3: Insufficient information to make a judgment.
Your answer should be only one number, such as 1, referring to the option.
A.6 Complete Results of Real-life Tasks
Daily plan result from GPT-4:
10:00 - 11:00 : Breakfast
11:00 - 12:00 : House Cleaning
12:00 - 13:00 : Free Time
13:00 - 15:00 : Basketball Playing
15:00 - 16:00 : Free Time
16:00 - 17:00 : Lunch
17:00 - 18:00 : Grocery Shopping
18:00 - 19:00 : Homework
19:00 - 20:00 : Laundry and Supper
20:00 - 23:00 : Continue Homework
In this plan, basketball playing is scheduled from 13.00 to 15.00 with an hour after breakfast to avoid playing right after eating as per the constraint. Lunch is then scheduled after basketball and Shopping done after that. Since laundry doesnt take time, it has been added with supper and rest of the time is allocated to homework.
Daily plan result after applying Formal-LLM:
10:00 - 10:00 Turning on the washer/laundry machine.
10:00 - 11:00 House cleaning.
11:00 - 12:00 Eating breakfast.
12:00 - 13:00 Grocery shopping.
13:00 - 15:00 Playing basketball.
15:00 - 16:00 Eating lunch.
16:00 - 19:00 Doing homework.
19:00 - 20:00 Eating supper.
Cooking recipe result from GPT-4:
Ingredients Preparation
1. Sliced Beef: 300 grams
2. Chinese Broccoli (Gai Lan): As needed, cut into segments
3. Scallions: 1 stalk, cut into sections
4. Ginger: As needed, sliced
5. Garlic: As needed, minced
6. Water: As needed
7. Cooking Oil: As needed
8. Salt: As needed
9. Light Soy Sauce: 2 tablespoons
10. Cooking Wine: 1 tablespoon
11. White Pepper Powder: As needed
12. Sugar: 1 tablespoon
13. Vinegar: 1 tablespoon
Cooking Utensils and Tableware Preparation
1. Stir-fry Pan: 1
2. Cooking Pot: 1
3. Tableware: Chopsticks, Soup Spoon, Stir-fry Pan Spoon, Several Bowls
Cooking Steps
Preparatory Steps
1. Marinate the beef slices with salt and cooking wine, stirring evenly, and let it marinate for 10 minutes.
2. Heat the stir-fry pan and add an appropriate amount of cooking oil.
Sautéing the Beef
1. Once the oil is hot, quickly stir-fry the marinated beef slices until they change color.
2. Add scallion sections, ginger slices, and minced garlic. Continue stirring until a fragrant aroma is released.
Sautéing the Chinese Broccoli
1. In the stir-fry pan, add water. Once the water boils, add the Chinese broccoli.
2. Quickly stir-fry the Chinese broccoli until it changes color, and season with salt.
Seasoning
1. Add light soy sauce, white pepper powder, sugar, and vinegar. Stir well.
2. Continue stirring to ensure that the seasoning is evenly coated on the beef and Chinese broccoli.
Completion
1. Ensure the beef is cooked through, and the Chinese broccoli is flavorful. Remove from heat.
Serving
1. Transfer the Chinese Broccoli Beef to bowls.
2. Serve with rice or noodles.
Tips
1. Avoid marinating the beef slices for too long to maintain a good texture.
2. Control the sautéing time for Chinese broccoli to preserve its crisp and tender texture.
3. Adjust the seasoning according to personal taste preferences, adding or reducing the amount of salt, sugar, and vinegar.
Cooking recipe result after applying Formal-LLM:
Here is the final plan:
Step 1: Wash raw beef slices with water.
Step 2: Prepare the seasoning: ginger, garlic, cooking oil, salt, light soy sauce, cooking wine, white pepper for the step: "Marinate clean slices of beef in a bowl with seasoning".
Step 3: Marinate clean slices of beef in a bowl with seasoning.
Step 4: Stir-fry the marinaded beef in a wok.
Step 5: Wash broccoli with water.
Step 6: Blanch clean broccoli in a cooking pot.
Step 7: Use a bowl to drain the blanched broccoli.
Step 8: Combine lightly cooked beef and blanched and drained broccoli in a wok.
Step 9: Prepare the seasoning: ginger, garlic, cooking oil, salt, light soy sauce, cooking wine, white pepper for the step: "Stir-fry the beef and broccoli mixture with the seasoning in a wok".
Step 10: Stir-fry the beef and broccoli mixture with the seasoning in a wok.
Step 11: Then, we get the cooked broccoli beef.
Carefully cooking at each step!
Risk management result from GPT-4:
Risk Management Plan:
Objective: To ensure a beneficial acquisition of Blizzard Entertainment by Microsoft while managing and mitigating potential risks.
I. Risk Identification
1. Financial Risks: A complete and thorough financial audit of Blizzard must be performed to determine potential financial risks, such bright as debts, pending lawsuit settlements, past or ongoing tax issues, declining revenue, etc.
2. Operational Risks: Understand Blizzards operational weaknesses and strengths, and the impact of the acquisition on Microsofts operations.
3. Legal Risks: Review Blizzards legal compliance status in all jurisdictions it operates in. Especially, risks related to intellectual property rights, privacy and data security laws, and labor laws must be evaluated.
4. Regulatory Risks: Determine any potential regulatory approvals Microsoft might need for the acquisition.
II. Risk Assessment:
1. Quantify each identified risk wherever possible, in terms of potential financial impact, time delay, brand damage etc.
2. Determine the likelihood of each risk occurrence.
3. Prioritize the risks based on their potential impact and the likelihood of occurrence.
III. Risk Mitigation Planning:
1. Financial Risk: This could be mitigated through leverage and insurance, renegotiation of terms, or such other financial instruments.
2. Operational Risk: To mitigate this risk, Microsoft can plan to gradually integrate Blizzard, keeping key management personnel, or finding synergies with Microsofts current operations.
3. Legal Risk: Microsoft should get warranties and indemnities from Blizzards owners to cover potential post-acquisition legal disputes.
4. Regulatory Risk: Early engagement with regulators and professional advice can help smooth the acquisition process.
IV. Risk Monitoring:
A system to continuously monitor and manage these risks should be set up. This could involve regular reporting, monitoring industry developments, legal changes etc.
V. Contingency Planning:
Establish contingency plans for high-impact risks. For instance, a backup plan if the acquisition fails, or if post-acquisition integration does not go as planned.
VI. Communication:
Continual communication with stakeholders to obtain feedback and to convey risk management efforts. This will ensure that all parties are aware of the risks and the measures being taken to mitigate them.
VII. Insurance Cover:
Consider adequate insurance coverage that includes Directors and Officers Liability and Professional Indemnity Insurance for all potential issues that might arise in connection with the proposed transaction.
The above steps will form the roadmap for Microsoft while negotiating terms for the acquisition of Blizzard Entertainment. Understanding risks before they come up ensures that Microsoft isnt surprised during negotiations, hence effectively mitigating them.
Risk management result after applying Formal-LLM:
Here is the final plan for Microsoft:
1: We need more information to ensure Shareholders of Blizzard Entertainment is entitled to sell Blizzard Entertainment. But we assume we have found out enough information to proceed the risk assessment process.
2: Then, based on current information, we believe Blizzard Entertainment is in good standing without material outstanding liabilities.
3: Then, based on current information, we believe neither CFIUS filings nor foreign investment filings in China are needed.
4: Then, we need to submit the regional antitrust filing of the United States and/or China. But we assume the filings will be approved to proceed the risk assessment process.
5: Pass the preliminary risk analysis. Details to be discussed in the transaction documents.
Note: This is a risk assessment process provided based on current information. Please ensure the accuracy of the provided information and possible additional supplemental information.
Appendix B Potential Risk
Our research focuses on making LLM-based agent planning more controllable. To the best of our knowledge, we do not notice the potential risk of our research.
Appendix C Additional information of Datasets
In our research, we used two benchmark datasets, OpenAGI (Ge et al., 2023a) and TravelPlanner (Xie et al., 2024). They are both under MIT license, which means these two datasets are publicly available.
OpenAGI (Ge et al., 2023a) is used as the agent creation package to build agents for LLM-based agents. TravelPlanner (Xie et al., 2024) is a benchmark crafted for evaluating language agents in tool-use and complex planning within multiple constraints. Our work is to explore the controllable planning ability of LLM-based agents, which is consistent with their intended use.
These two benchmark datasets are synthesized by algorithm and thus, do not contain any individual personal information.
The language of these two benchmark datasets is English.
For the number of examples and details of train / test / dev splits, we use the same dataset and follow the the same data splitting method with these two papers (Ge et al., 2023a; Xie et al., 2024).